### v3

@@ 4.2 Results of RQ1: Mined Architectural Information
Architectural technical debt: Architectural Technical Debt (ATD) is incurred by architecture decisions that consciously or unconsciously compromise system-wide QAs, particularly maintainability and evolvability (Li et al., 2014). Typical ATD includes violations of architecture design principles or rules (Li et al., 2015). ATD needs to be identified and removed since it is harmful to the system’s long-term health (Li et al., 2015). 24 out of 104 studies (23.0%) developed approaches for detecting and mining ATD for different purposes, such as mining ATD for further management in order to keep the accumulated ATD under control [S37]. We categorized the mined architectural information related to ATD into three subcategories, namely architectural smell (e.g., Unstable Interface, Modularity Violations, Hublike Dependency, Chatty Service, Tiny Service), architectural anti-pattern (e.g., Clique, Crossing), and architectural compliance issue (see Table 5). For instance, in [S71], Díaz-Pace et al. developed an approach based on Link Prediction (LP) techniques and an ML classification model to mine two types of architectural smells, namely cyclic dependency and hub-like dependency from OSS projects. On the other hand, Mo et al. [S64] presented an approach for automatically detecting and mining six types of architecture anti-patterns (including unstable interface, modularity violation groups), defined as connections among files that violate design principles. This mined architectural information can be used by architects to pinpoint architectural anti-patterns in the systems, quantify their severity, and determine priorities for refactoring. Maffort et al. [S34] proposed an approach that relies on four heuristic models for detecting and mining architectural compliance issues. Specifically, the approach mines the absences (i.e., something expected is not found) and divergences (i.e., something prohibited is found) in source code based architectures. The mined architectural compliance issues can be used to rapidly raise architectural deviation warnings, without deeply involving architects.
@@
Architectural change: Software changes are inevitable. There are many reasons for software changes, such as repairing defects and evolving user requirements (Williams and Carver, 2010). When changes affect the architecture, software engineers need a comprehensive understanding of the causes of architecture changes and their impact (Williams and Carver, 2010). This understanding is important because, as changes may affect many aspects of a system and introduce complexities in architecture, which will likely introduce a significant number of architectural issues. 18 out of 104 studies (17.3%) analyzed architectural change information from different perspectives (e.g., classification of architectural changes [S40]) to support the development. We adopted a predefined classification of architectural changes introduced by Williams and Carver (Williams and Carver, 2010) to categorize the mined architectural change into four subcategories: perfective — indicates new requirements and improved functionality, corrective — addresses flaws, adaptive — occurs for new environment or for imposing new policies, and preventative — indicates restructuring or redesigning the system. For example, Mondal [S40] utilized an ML-based approach for mining information related to architectural changes from source code and commit messages, and labeled them into four subcategories, namely perfective change, corrective change, adaptive change, and preventive change. This mined architectural information can help developers better analyze and characterize the causes and impact of architectural changes prior to their implementations. Figure 6(b) shows an example of mined architectural change (i.e., perfective change) from the developer mailing list.
@@
Design relationship: An architecture of a system is designed with the aid of architectural elements (e.g., tactics and components). However, there are complex interdependency relationships between these elements, and many of the tasks faced by developers today are to deal with efficient and effective use of these elements together during the application design to meet both functional and non-functional goals (Bi et al., 2018). For example, designing with a certain architecture pattern to address specific quality attribute requirements cannot be considered in isolation, but there might be other key architectural elements to be considered, such as design contexts (e.g., application context) (Bedjeti et al., 2017). Therefore, 17 out of 104 (16.3%) studies researched on mining design relationships from various sources to aid the development. We categorized the mined architectural information in the Design relationship category into seven subcategories, such as Component-Component relationship, Architectural tactic-Context relationship (see Table 5). For instance, Karthik et al. [S20] presented an automatic ML-based approach for mining the relationships between architectural components from the unstructured text of Q&A site posts. Moreover, the mined Component-Component relationships were categorized into three types, namely incompatibility, required, and recommended. This mined architectural information can help developers who work with component-based systems to effectively understand the Component-Component relationships of certain applications similar to the applications discussed on Q&A sites. Whilst applying Architectural Tactics (ATs) to address QAs is well explored in existing works, e.g., (Mirakhorli et al., 2013; Bi et al., 2021), there are no guidelines for architects, who look for information on what considerations (e.g., design contexts) they need to consider when applying ATs to address QA concerns. In order to provide architects with such information, Gopalakrishnan et al. [S27] proposed an ML and text mining-based approach for extracting and mining typical design contexts in the source code in which architectural tactics are implemented, and analyzed the relationship between the design contexts and architectural tactics implemented in the code (i.e., Architectural tactic-Context relationship). On the other hand, Wijerathna et al. [S87] employed an automatic ML-based approach for mining design contexts such as platform context (e.g., software context) from Stack Overflow posts, and manually structured the design relationships between design contexts and architectural patterns used in practice (i.e., Architectural pattern-Context relationship). This information can assist software engineers (e.g., architects) in choosing the appropriate architectural patterns with respect to design context. In [S31], Velasco-Elizondo et al. used an approach based on an information extraction technique (i.e., entity extraction) and knowledge representation (i.e., ontology) to automatically analyze and mine architecture patterns considering specific QAs (i.e., Architectural pattern-Quality attribute relationship) from architectural pattern descriptions. The mined architectural information in this subcategory can facilitate developers to select architectural patterns through knowing whether specific QAs are promoted or inhibited.
Finding 1: 7 main categories and 29 subcategories of architectural information are mined from various sources to assist architecting activities. Finding 2: Architectural description (69.2%, 72 out of 104 studies), architectural solution (42.3%, 44 out of 104 studies), and architectural decision (30.8%, 32 out of 104 studies) are the top three most mined categories of architectural information.
@@ 4.3 Results of RQ2: Sources used for Mining Architectural Information
Finding 3: VCS, e.g., GitHub, (used by 52 studies), Online Q&A sites, e.g., Stack Overflow, (used by 33 studies), and Wiki (used by 21 studies) are the top three most frequently used sources for mining architectural information. Finding 4: Five types of datasets are used for mining architectural information, in which code-based datasets are the most frequently employed datasets (52 out of 104 studies, 50%). Moreover, image-based datasets are rarely used and one reason could be that there is a lack of open source image-based datasets in software repositories.
@@ 4.4 Results of RQ3: Supported Architecting Activities
We collected the architecting activities that can be supported by the mined architectural information (the results of RQ1). As mentioned in Section 3.2.5, a hybrid approach (which is a combination of open coding & constant comparison, predefined classifications of architecting activities by Hofmeister et al. (2007), Tang et al. (2010), and Li et al. (2013), and descriptive statistics) was used for analyzing the data item D7 to answer RQ3. Hofmeister et al. (2007) presented a general model for architecture design, which consists of three main activities: architecture analysis, architecture synthesis, and architecture evaluation. Tang et al. (2010) extended this general model by adding two architecting activities (i.e., architecture implementation and architecture maintenance) in the architecture life cycle. These two architecting activities emphasize that architecture is not only an important asset for design but also for the later stage in the software development life cycle. Li et al. (2013) extended the aforementioned architecting activities by adding five more activities (i.e., architecture recovery, architectural description, architecture understanding, architecture impact analysis, and architecture reuse) from the perspective of Architecture Knowledge (AK) management (e.g., AK reuse) to support the architecture life cycle. Therefore, we followed the ten predefined architecting activities when answering RQ3. However, through our qualitative data analysis (see Section 3.2.5), we obtained one additional architecting activity, namely “architecture conformance checking”.
@@
Moreover, to comprehensively understand how the identified architecting activities are supported, we analyzed the primary studies to explore how these studies contribute to architecting activities through the proposed approaches and tools (results of RQ4). To achieve this, we investigated these primary studies to present the relationships between architecting activities (e.g., architecture reuse, architecture analysis) with respect to types of problems solved (e.g., architectural change classification), approaches and/or tools used (e.g., clustering-based approaches), mined architectural information (e.g., architectural views), dataset types (e.g., code-based, metric-based datasets) as well as sources (e.g., Code search engines) used in mining architectural information.
@@
We present the frequency of occurrence of the eleven architecting activities supported by the mined architectural information in the Architecting Activities sheet in the dataset (de Dieu et al., 2023). Moreover, that sheet also provides the relationships between supported architecting activities, mined architectural information, datasets, and sources used for mining architectural information as well as the relevant studies. These relationships act as a panorama to comprehensively understand the state of the research in mining architectural information. Furthermore, Figure 9 presents the results in a map through the studies distributed over the dimensions of architecting activities, mining approaches, and time period. In the top of Figure 9, we illustrate eleven architecting activities and three categories of mining approaches. The relationships of the three elements are shown in the lower half part of the map. The left part reflects the relationship between publication year and architecting activities of the selected studies, and the number in bubbles represents the corresponding selected studies (listed in Appendix A) in specific architecting activities. The right part shows the relationship between architecting activities and mining approaches. The number in bubbles signifies the selected studies employing a certain type of mining approach in a specific architecting activity. As shown in Figure 9, most of the studies treated the problems in architecting activities as classification and clustering tasks and, therefore, these studies mostly employed classification-based and clustering-based approaches and/or tools for mining architectural information from diverse sources in order to support various architecting activities.
@@
Architecture Understanding (AU) is conducted to comprehend elements of an architecture design, such as architectural solutions, architectural components, and relationships between those elements (e.g., the relationships between architectural components (Li et al., 2013; Stevanetic and Zdun, 2014)). The information gained during AU can be used as input for other architecting activities (such as architecture implementation) (Bengtsson et al., 2004). The AU activity is the most supported activity (i.e., 44.2% , 46 out of 104) of the studies (see the Architecting Activity sheet in the dataset (de Dieu et al., 2023)).
@@
Problems solved and approaches used in AU - Various problems, such as poor understanding of architectural elements (e.g., architectural patterns), poor understating of the relationships between architectural elements are faced by software engineers during architecture design (Li et al., 2013). One common practice of mining approaches is to process textual or non-textual artifacts and mine key information to improve the understanding of those artifacts, especially for large collections of textual or non-textual artifacts (Feldman and Sanger, 2007). The majority of the studies in AU activity treated the problems in this activity as classification and clustering-based tasks (see Figure 9)). Thus, these studies mainly utilized classification-based (e.g., [S19][S21][S37][S40][S69]) and clustering-based (e.g., [S67][S46][S74][S91]) approaches and tools for mining architectural information in order to solve different problems in this activity (see the Architecting Activities sheet in the dataset (de Dieu et al., 2023). We classified the problems solved in AU activity into seven types: understanding of architecture description (e.g., via mining architectural views [S91] and models [S39]), understanding of architectural solutions (e.g., via classification of architectural solutions [S61][S21]), understanding of architectural decisions (e.g., via classification of architectural decisions [S19]), understanding of system requirements (e.g., classification of system requirements [S83]), understating of architectural changes (e.g., via classification of architectural changes [S40][S77]), understanding of architectural technical debt (e.g., via detection and classification architectural smells [S29][S96][S37]), and understanding of the relationships between architectural elements (e.g., via mining Component-Component relationship (e.g., [S67][S74]). For instance, for improving architectural description via mining architectural views, Nascimento Vale et al. [S59] proposed an approach called Keecle for semi-automatically identification of architecturally relevant classes that represent components from source code. The approach uses Naïve Bayes classification algorithm to classify relevant classes that provide a high-level overview of a software system in order to help developers understand the entire software system.
@@
Datasets and sources used in AU - Other critical problems in AU activity are the datasets and sources in which architectural information can be extracted in order to solve the identified problems. Thus, we also explored and reported datasets and sources used in AU. Four types of datasets were used in AU: code-based (e.g., [S67][S74]), text-based (e.g., [S19][S58]), metric-based (e.g., [S46]), and mixed datasets (e.g., code and documentation [S69]) (see the Architecting Activities sheet in the dataset (de Dieu et al., 2023)). Moreover, these datasets were mainly collected from code repositories, such as GitHub (e.g., [S47][S37]).
@@
Architecture Maintenance and Evolution (AME): No software product is ever completely finished and all need some form of ongoing maintenance. Architectural maintenance is to adjust the architecture according to changes and faults (Tang et al., 2010). Architecture may evolve during its lifetime to respond to changing requirements (Tang et al., 2010). AME is the second most assisted architecting activity (i.e., 40.4%, 42 out of 104 studies) (see the Architecting Activity sheet in the dataset (de Dieu et al., 2023)).
@@
Datasets and sources used in AME - Common datasets in AME are: code-based (e.g., [S37][S90][S95]), metric-based (e.g., [S46]), and mixed (e.g., code and commit messages) dataset (e.g., [S40]), and these datasets were mainly collected from version control systems.
Architecture Recovery (ARec) is mainly performed to recover architectural information from systems implementation (Riva, 2000) or documentation (Shahin et al., 2013). The ARec activity is the third most supported activity (i.e., 32.7% (34 out of 104)) of the studies.
Problems solved and approaches used in ARec - Several problems have been solved in ARec activity including recovering architectural description (e.g., via mining architectural views [S46]), recovering architectural solutions (e.g., via mining architectural patterns [S1]), recovering architectural decisions (e.g., via extraction of architectural decisions [S19]), recovering system requirements (e.g., via extraction of system requirements [S76]) (see the Architecting Activities sheet in the dataset (de Dieu et al., 2023)). Moreover, the studies in ARec treated the problems in this activity as classification, clustering, and Information Retrieval (IR) based tasks and used classification-based (e.g., [S19][S33]), clustering-based (e.g., [S41][S46][S47]), and IR-based (e.g., [S28][S53]) approaches and/or tools for mining architectural information in order to support ARec. As shown in Figure 9, clustering-based and classification-based approaches are the most used approaches in ARec activity. For example, for clustering-based approaches, Shahbazian et al. [S47] developed an approach named RecovAr that relies on two state-of-the-art architectural recovery techniques (i.e., Algorithm for Comprehension-Driven Clustering (ACDC) (Tzerpos and Holt, 2000) and Architecture Recovery using Concerns (ARC) (Garcia et al., 2011)) for automatically recovering architectural decisions (e.g., behavioral decisions) from source code of software projects.
@@
Architecture Implementation (AIm): During the AIm activity, developers transform and implement architecture design into code (Tang et al., 2010). The AIm activity is supported by 19.2% (20 out of 104) of the studies (see the Architecting Activities sheet in the dataset (de Dieu et al., 2023)).
@@
Architecture Analysis (AA): AA examines, filters, and reformulates architectural concerns and context that an architecture needs to address (Hofmeister et al., 2007). During AA, Architecturally Significant Requirements (ASRs) are identified from given architectural concerns and context (Hofmeister et al., 2007). The AA activity is supported by 18.3% (19 out of 104) of the studies (see the Architecting Activities sheet in the dataset (de Dieu et al., 2023)).
@@
Datasets and sources AReu - Code-based (e.g., [S13]), text-based (e.g., [S1][S25]), and mixed (e.g., code and documentation [S69]) datasets are the three types of datasets used in AReu activity. On the other hand, version control systems, Q&A sites, Wiki, issue tracking systems, and technical blogs and tutorials are the prevalent sources for AReu (see the Architecting Activities sheet in the dataset (de Dieu et al., 2023)).
Architecture Evaluation (AE) aims to assess the architectural solutions that are proposed during AS activity against the ASRs (Hofmeister et al., 2007). During the AE activity, various factors are considered, such as pros and cons, trade-offs, and constraints of architectural solutions that result in the selection of appropriate architectural solutions for satisfying ASRs. The AE activity is supported by 17.3% (18 out of 104) of the studies (see the Architecting Activities sheet in the dataset (de Dieu et al., 2023)).
@@
Datasets and sources in AE - Code-based (e.g., [S49]) and text-based (e.g., [S2][S14]) datasets are the common datasets in AE activity. On the other hand, Q&A sites, Wiki, issue tracking systems, and technical blogs and tutorials are the most employed sources in this activity.

### v4

@@ 4.2 Results of RQ1: Mined Architectural Information
Architectural technical debt: Architectural Technical Debt (ATD) is incurred by architecture decisions that consciously or unconsciously compromise system-wide QAs, particularly maintainability and evolvability (Li et al., 2014). Typical ATD includes violations of architecture design principles or rules (Li et al., 2015). ATD needs to be identified and removed since it is harmful to the system’s long-term health (Li et al., 2015). 24 out of 104 studies (23.0%) developed approaches for detecting and mining ATD for different purposes, such as mining ATD for further management in order to keep the accumulated ATD under control [S37]. We categorized the mined architectural information related to ATD into three subcategories, namely architectural smell (e.g., Unstable Interface, Modularity Violations, Hublike Dependency, Chatty Service, Tiny Service), architectural anti-pattern (e.g., Clique, Crossing), and architectural compliance issue (see Table LABEL:minedArchitecturalInfo). For instance, in [S71], Díaz-Pace et al. developed an approach based on Link Prediction (LP) techniques and an ML classification model to mine two types of architectural smells, namely cyclic dependency and hub-like dependency from OSS projects. On the other hand, Mo et al. [S64] presented an approach for automatically detecting and mining six types of architecture anti-patterns (including unstable interface, modularity violation groups), defined as connections among files that violate design principles. This mined architectural information can be used by architects to pinpoint architectural anti-patterns in the systems, quantify their severity, and determine priorities for refactoring. Maffort et al. [S34] proposed an approach that relies on four heuristic models for detecting and mining architectural compliance issues. Specifically, the approach mines the absences (i.e., something expected is not found) and divergences (i.e., something prohibited is found) in source code based architectures. The mined architectural compliance issues can be used to rapidly raise architectural deviation warnings, without deeply involving architects.
@@
Architectural changes: Software changes are inevitable. There are many reasons for software changes, such as repairing defects and evolving user requirements (Williams and Carver, 2010). When changes affect the architecture, software engineers need a comprehensive understanding of the causes of architecture changes and their impact (Williams and Carver, 2010). This understanding is important because, as changes may affect many aspects of a system and introduce complexities in architecture, which will likely introduce a significant number of architectural issues. 18 out of 104 studies (17.3%) analyzed architectural change information from different perspectives (e.g., classification of architectural changes [S40]) to support the development. We adopted a predefined classification of architectural changes introduced by Williams and Carver (Williams and Carver, 2010) to categorize the mined architectural change into four subcategories: perfective — indicates new requirements and improved functionality, corrective — addresses flaws, adaptive — occurs for new environment or for imposing new policies, and preventative — indicates restructuring or redesigning the system. For example, Mondal [S40] utilized an ML-based approach for mining information related to architectural changes from source code and commit messages, and labeled them into four subcategories, namely perfective change, corrective change, adaptive change, and preventive change. This mined architectural information can help developers better analyze and characterize the causes and impact of architectural changes prior to their implementations. Figure 6(b) shows an example of mined architectural change (i.e., perfective change) from the developer mailing list.
@@
Design relationships: An architecture of a system is designed with the aid of architectural elements (e.g., tactics and components). However, there are complex interdependency relationships between these elements, and many of the tasks faced by developers today are to deal with efficient and effective use of these elements together during the application design to meet both functional and non-functional goals (Bi et al., 2018). For example, designing with a certain architecture pattern to address specific quality attribute requirements cannot be considered in isolation, but there might be other key architectural elements to be considered, such as design contexts (e.g., application context) (Bedjeti et al., 2017). Therefore, 17 out of 104 (16.3%) studies researched on mining design relationships from various sources to aid the development. We categorized the mined architectural information in the Design relationship category into seven subcategories, such as Component-Component relationship, Architectural tactic-Context relationship (see Table LABEL:minedArchitecturalInfo). For instance, Karthik et al. [S20] presented an automatic ML-based approach for mining the relationships between architectural components from the unstructured text of Q&A site posts. Moreover, the mined Component-Component relationships were categorized into three types, namely incompatibility, required, and recommended. This mined architectural information can help developers who work with component-based systems to effectively understand the Component-Component relationships of certain applications similar to the applications discussed on Q&A sites. Whilst applying Architectural Tactics (ATs) to address QAs is well explored in existing works, e.g., (Mirakhorli et al., 2013; Bi et al., 2021), there are no guidelines for architects, who look for information on what considerations (e.g., design contexts) they need to consider when applying ATs to address QA concerns. In order to provide architects with such information, Gopalakrishnan et al. [S27] proposed an ML and text mining-based approach for extracting and mining typical design contexts in the source code in which architectural tactics are implemented, and analyzed the relationship between the design contexts and architectural tactics implemented in the code (i.e., Architectural tactic-Context relationship). On the other hand, Wijerathna et al. [S87] employed an automatic ML-based approach for mining design contexts such as platform context (e.g., software context) from Stack Overflow posts, and manually structured the design relationships between design contexts and architectural patterns used in practice (i.e., Architectural pattern-Context relationship). This information can assist software engineers (e.g., architects) in choosing the appropriate architectural patterns with respect to design context. In [S31], Velasco-Elizondo et al. used an approach based on an information extraction technique (i.e., entity extraction) and knowledge representation (i.e., ontology) to automatically analyze and mine architecture patterns considering specific QAs (i.e., Architectural pattern-Quality attribute relationship) from architectural pattern descriptions. The mined architectural information in this subcategory can facilitate developers to select architectural patterns through knowing whether specific QAs are promoted or inhibited.
Category Subcategory Studies Count
Architectural description (69.2%, 72 out of 104) Architectural model [S11] [S12] [S29] [S30] [S36] [S37] [S39] [S46] [S49] [S51] [S54] [S62] [S64] [S67] [S68] [S73] [S75] [S80] [S83] [S86] 20
Architectural view [S8] [S16] [S18] [S30] [S35] [S37] [S39] [S46] [S47] [S54] [S59] [S61] [S64] [S68] [S98] [S70] [S81] [S102] 18
Architectural rationale [S1] [S2] [S4] [S11] [S18] [S23] [S24] [S25] [S26] [S28] [S38] [S47] [S79] 13
System of interest [S11] [S26] [S29] [S39] [S60] [S72] [S85] [S100] [S103] 9
Architectural concern [S4] [S6] [S7] [S8] [S9] [S11] [S38] [S66] 8
Architectural solution (42.3%, 44 out of 104) Architectural pattern [S1] [S2] [S3] [S4] [S16] [S23] [S24] [S28] [S32] [S44] [S49] [S50] [S52] [S53] [S54] [S61] [S63] [S69] [S70] [S82] [S79] 21
Architectural tactic [S1] [S4] [S6] [S7] [S13] [S26] [S9] [S24] [S49] [S50] [S53] [S56] [S57] [S79] 14
Framework [S3] [S4] [S14] [S16] [S26] [S49] [S50] [S52] [S61] 9
Architectural decision (30.8%, 32 out of 104) Technology decision [S2] [S3] [S4] [S14] [S15] [S16] [S18] [S23] [S33] [S38] [S53] [S85] 12
Structural decision [S17] [S18] [S19] [S22] [S23] [S32] [S33] [S47] [S54] [S85] 10
Behavioral decision [S18] [S19] [S22] [S33] [S42] [S47] [S54] 7
Ban or non-existence decision [S19] [S22] [S33] 3
System requirement (29.8%, 31 out of 104) Quality attribute [S1] [S2] [S3] [S4] [S8] [S11] [S16] [S17] [S23] [S24] [S26] [S28] [S32] [S42] [S61] [S66] [S76] [S83] 18
Functional requirement [S1] [S4] [S17] [S43] [S45] [S54] [S55] [S83] [S79] 9
Constraint [S1] [S2] [S3] [S54] 4
Architectural technical debt (23.0%, 24 out of 104) Architectural smell [S5] [S29] [S65] [S71] [S84] [S88] [S90] [S92] [S93] [S94] [S97] [S98] [S99] [S100] 14
Architectural anti-pattern [S37] [S60] [S64] [S72] [S95] [S104] 6
Architectural compliance issue [S30] [S34] [S36] [S73] 4
Architectural change (17.3%, 18 out of 104) Perfective change [S17] [S18] [S32] [S35] [S40] [S11] [S77] 7
Corrective change [S35] [S32] [S40] [S46] [S77] 5
Adaptive change [S35] [S40] [S77] 3
Preventative change [S35] [S40] [S77] 3
Design relationship (16.3%, 17 out of 104) Component-Component relationship [S12] [S20] [S48] [S59] [S67] [S69] [S74] [S101] [S103] 9
Architectural tactic-Context relationship [S6] [S27] 2
Architectural pattern-Architectural pattern relationship [S63] [S78] 2
Architectural pattern-Context relationship [S87] 1
Architectural pattern-Quality attribute relationship [S31] 1
Architectural tactic-Quality attribute relationship [S6] 1
Design pattern-Architectural tactic relationship [S21] 1
Key Findings of RQ1 Finding 1: A broad range of architectural information has been mined to support the development. The mined architectural information, such as Architectural tactic-Context relationship, can assist software engineers (e.g., developers) in choosing and implementing the appropriate architectural tactics with respect to design context.
@@ 4.3 Results of RQ2: Sources used for Mining Architectural Information
Key Findings of RQ2 Finding 2: The majority of primary studies utilized code-based and text-based datasets for mining architectural information. Moreover, these datasets were mainly collected from VCS, Online Q&A sites, and Wiki. One possible reason is that code-based and text-based datasets are usually the major datasets available in open-source repositories. Furthermore, image-based datasets are rarely used and one reason could be that there is a lack of open source image-based datasets in software repositories.
@@ 4.4 Results of RQ3: Supported Architecting Activities
As mentioned in Section 3.2.5, a hybrid approach (which is a combination of open coding & constant comparison, predefined classifications of architecting activities by Hofmeister et al. (2007), Tang et al. (2010), and Li et al. (2013), and descriptive statistics) was used for analyzing the data item D7 to answer RQ3. Hofmeister et al. (2007) presented a general model for architecture design, which consists of three main activities: architecture analysis, architecture synthesis, and architecture evaluation. Tang et al. (2010) extended this general model by adding two architecting activities (i.e., architecture implementation and architecture maintenance) in the architecture life cycle. These two architecting activities emphasize that architecture is not only an important asset for design but also for the later stage in the software development life cycle. Li et al. (2013) extended the aforementioned architecting activities by adding five more activities (i.e., architecture recovery, architectural description, architecture understanding, architecture impact analysis, and architecture reuse) from the perspective of Architecture Knowledge (AK) management (e.g., AK reuse) to support the architecture life cycle. Therefore, we followed the ten predefined architecting activities when answering RQ3. However, through our qualitative data analysis (see Section 3.2.5), we obtained one additional architecting activity, namely “architecture conformance checking”.
@@
Moreover, to comprehensively understand how the identified architecting activities are supported, we analyzed the primary studies to explore how these studies contribute to architecting activities through the proposed approaches and tools (results of RQ4). To achieve this, we investigated these primary studies to present the relationships between architecting activities (e.g., architecture reuse) with respect to types of problems solved (e.g., architectural change classification), approaches and/or tools used (e.g., clustering-based approaches), mined architectural information (e.g., architectural views), dataset types (e.g., code-based datasets) as well as sources (e.g., Code search engines) used in mining architectural information.
@@
We present the frequency of occurrence of the eleven architecting activities supported by the mined architectural information in the Architecting Activities sheet in the Supplementary Material (de Dieu et al., 2024). Moreover, that sheet also provides the relationships between supported architecting activities, mined architectural information, datasets, and sources used for mining architectural information as well as the relevant studies. These relationships act as a panorama to comprehensively understand the state of the research in mining architectural information. Furthermore, Figure 9 presents the results in a map through the studies distributed over the dimensions of architecting activities, mining approaches, and time period. In the top of Figure 9, we illustrate eleven architecting activities and three categories of mining approaches. The relationships of the three elements are shown in the lower half part of the map. The left part reflects the relationship between publication year and architecting activities of the selected studies, and the number in bubbles represents the corresponding selected studies (see the Selected Studies in the Supplementary Material (de Dieu et al., 2024)) in specific architecting activities. The right part shows the relationship between architecting activities and mining approaches. The number in bubbles signifies the selected studies employing a certain type of mining approach in a specific architecting activity. As shown in Figure 9, most of the studies treated the problems in architecting activities as classification and clustering tasks and, therefore, these studies mostly employed classification-based and clustering-based approaches and/or tools for mining architectural information from diverse sources to support various architecting activities.
@@
Architecture Understanding (AU) is conducted to comprehend elements of an architecture design, such as architectural solutions, architectural components, and relationships between those elements (e.g., the relationships between architectural components (Stevanetic and Zdun, 2014)). The information gained during AU can be used as input for other architecting activities (such as architecture implementation). The AU activity is the most supported activity (i.e., 44.2% , 46 out of 104) of the studies (see the Architecting Activity sheet in the Supplementary Material (de Dieu et al., 2024)).
@@
Problems solved and approaches used in AU - Various problems, such as poor understanding of architectural elements (e.g., architectural patterns), poor understating of the relationships between architectural elements are faced by software engineers during architecture design (Li et al., 2013). One common practice of mining approaches is to process textual or non-textual artifacts and mine key information to improve the understanding of those artifacts, especially for large collections of textual or non-textual artifacts. The majority of the studies in AU activity treated the problems in this activity as classification and clustering-based tasks (see Figure 9)). Thus, these studies mainly utilized classification-based (e.g., [S19][S21][S37][S40][S69]) and clustering-based (e.g., [S67][S46][S74][S91]) approaches and tools for mining architectural information in order to solve different problems in this activity (see the Architecting Activities sheet in the Supplementary Material (de Dieu et al., 2024)). We classified the problems solved in AU activity into seven types: understanding of architecture description (e.g., via mining architectural views [S91] and models [S39]), understanding of architectural solutions (e.g., via classification of architectural solutions [S61][S21]), understanding of architectural decisions (e.g., via classification of architectural decisions [S19]), understanding of system requirements (e.g., classification of system requirements [S83]), understating of architectural changes (e.g., via classification of architectural changes [S40][S77]), understanding of architectural technical debt (e.g., via detection and classification architectural smells [S29][S96][S37]), and understanding of the relationships between architectural elements (e.g., via mining Component-Component relationship (e.g., [S67][S74]). For instance, for improving architectural description via mining architectural views, Nascimento Vale et al. [S59] proposed an approach called Keecle for semi-automatically identification of architecturally relevant classes that represent components from source code. The approach uses Naïve Bayes classification algorithm to classify relevant classes that provide a high-level overview of a software system in order to help developers understand the entire software system.
@@
Datasets and sources used in AU - Other critical problems in AU activity are the datasets and sources in which architectural information can be extracted in order to solve the identified problems. Thus, we also explored and reported datasets and sources used in AU. Four types of datasets were used in AU: code-based (e.g., [S67][S74]), text-based (e.g., [S19][S58]), metric-based (e.g., [S46]), and mixed datasets (e.g., code and documentation [S69]) (see the Architecting Activities sheet in the Supplementary Material (de Dieu et al., 2024)). Moreover, these datasets were mainly collected from code repositories, such as GitHub (e.g., [S47][S37]).
@@
Architecture Maintenance and Evolution (AME): No software product is ever completely finished and all need some form of ongoing maintenance. Architectural maintenance is to adjust the architecture according to changes and faults. Architecture may evolve during its lifetime to respond to changing requirements (Tang et al., 2010). AME is the second most assisted architecting activity (i.e., 40.4%, 42 out of 104 studies) (see the Architecting Activity sheet in the Supplementary Material (de Dieu et al., 2024)).
@@
Datasets and sources used in AME - Common datasets in AME are: code-based (e.g., [S37][S90][S95]), metric-based (e.g., [S46]), and mixed (e.g., code and commit messages) (e.g., [S40]) datasets, and these datasets were mainly collected from version control systems.
Architecture Recovery (ARec) is mainly performed to recover architectural information from systems implementation Schmitt Laser et al. (2020) or documentation (Shahin et al., 2013). The ARec activity is the third most supported activity (i.e., 32.7% (34 out of 104)) of the studies.
Problems solved and approaches used in ARec - Several problems have been solved in ARec activity including recovering architectural description (e.g., via mining architectural views [S46]), recovering architectural solutions (e.g., via mining architectural patterns [S1]), recovering architectural decisions (e.g., via extraction of architectural decisions [S19]), recovering system requirements (e.g., via extraction of system requirements [S76]) (see the Architecting Activities sheet in the Supplementary Material (de Dieu et al., 2024)). Moreover, the studies in ARec treated the problems in this activity as classification, clustering, and Information Retrieval (IR) based tasks and used classification-based (e.g., [S19][S33]), clustering-based (e.g., [S41][S46][S47]), and IR-based (e.g., [S28][S53]) approaches and/or tools for mining architectural information in order to support ARec. As shown in Figure 9, clustering-based and classification-based approaches are the most used approaches in ARec activity. For example, for clustering-based approaches, Shahbazian et al. [S47] developed an approach named RecovAr that relies on two state-of-the-art architectural recovery techniques (i.e., Algorithm for Comprehension-Driven Clustering (ACDC) (Tzerpos and Holt, 2000) and Architecture Recovery using Concerns (ARC) [S41]) for automatically recovering architectural decisions (e.g., behavioral decisions) from source code of software projects.
@@
Architecture Implementation (AIm): During the AIm activity, developers transform and implement architecture design into code (Hofmeister et al., 2007). The AIm activity is supported by 19.2% (20 out of 104) of the studies (see the Architecting Activities sheet in the Supplementary Material (de Dieu et al., 2024)).
@@
Architecture Analysis (AA): AA examines, filters, and reformulates architectural concerns and context that an architecture needs to address (Li et al., 2013). During AA, Architecturally Significant Requirements (ASRs) are identified from given architectural concerns and context (Hofmeister et al., 2007). The AA activity is supported by 17.3% (18 out of 104) of the studies (see the Architecting Activities sheet in the Supplementary Material (de Dieu et al., 2024)).
@@
Datasets and sources AReu - Code-based (e.g., [S13]), text-based (e.g., [S1][S25]), and mixed (e.g., code and documentation [S69]) datasets are the three types of datasets used in AReu activity. On the other hand, version control systems, Q&A sites, Wiki, issue tracking systems, and technical blogs and tutorials are the prevalent sources for AReu (see the Architecting Activities sheet in the Supplementary Material (de Dieu et al., 2024)).
Architecture Evaluation (AE) aims to assess the architectural solutions that are proposed during AS activity against the ASRs (Hofmeister et al., 2007). During the AE activity, various factors are considered, such as pros and cons, trade-offs, and constraints of architectural solutions that result in the selection of appropriate architectural solutions for satisfying ASRs. The AE activity is supported by 16.3% (17 out of 104) of the studies (see the Architecting Activities sheet in the Supplementary Material (de Dieu et al., 2024)).
@@
Datasets and sources in AE - Code-based (e.g., [S34]) and text-based (e.g., [S2][S14]) datasets are the common datasets in AE activity. On the other hand, Q&A sites, Wiki, issue tracking systems, and technical blogs and tutorials are the most employed sources in this activity.